AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 22 of 33  ELOG logo
ID Date Author Subject
  244   Tue Feb 22 21:09:39 2022 David, YuriCurrent status

Pressure seems as before 19:10.

Merger, Tape server (WS3) checks: DONE, writing ~1 MB/s, refreshing is fine

Leakage current (WS4): 3.47 uA , stable current

DAQ (WS2): temperature checked, clock status:ok, adc calibration:ok, white rabbit check:ok

DAQ (WS2): everything ok

Online monitoring (WS 5): done

  243   Tue Feb 22 20:08:40 2022 David, YuriCurrent status

Recently the pressure started to raise a bit, we have increased the temperature from 129.1 K to 129.2 K. Current pressure at S1 2.10e-7.

Merger, Tape server (WS3) checks: DONE, writing ~1 MB/s, refreshing is fine

Leakage current (WS4): 3.46 uA , stable current

DAQ (WS2): temperature checked, clock status:ok, adc calibration:ok, white rabbit check:ok

DAQ (WS2): everything ok

Online monitoring (WS 5): done

  242   Tue Feb 22 19:28:35 2022 David, Yurivacuum

At 19:10 Frank called us that vacuum in sections 08 and 10 got worse by about 1.5 times. There is a possible danger that vacuum valves can get closed if rest gas pressure is above the set treshhold.

Nikos warmed the cryostat by 0.5 degrees from 129.0 K to 129.5 K and the pressure became better than before. We cooled the system in steps of 0.1 degress down to 129.1 K and again observed the pressure rise.

We think we can stay with the present conditions. However, the pressure shall be monitored not to hit the treshhold.

  241   Tue Feb 22 19:11:18 2022 DavidCurrent status

Still trying to recover previous pressures, more details in the next entry.

Merger, Tape server (WS3) checks: DONE, writing ~1030 kB/s, refreshing is fine

Leakage current (WS4): 3.46 uA , stable current

DAQ (WS2): temperature checked,

DAQ (WS2): everything ok

Online monitoring (WS 5): done;

  240   Tue Feb 22 18:23:02 2022 David. YuryVacuum problem

At 19:10 the vacuum has increased suddenly everywhere in the ring, roughly a factor 1.5.

  239   Tue Feb 22 17:32:38 2022 DavidCurrent status

Michael is modifying the settings of the electron cooler --> changed beam from bunched to coasting

Merger, Tape server (WS3) checks: DONE, writing ~1036 kB/s, refreshing is fine

Leakage current (WS4): 3.46 uA , stable current

DAQ (WS2): temperature checked,

     - clock status:

Clock status test result: Passed 5, Failed 0

Understand status as follows
Status bit 3 : firmware PLL that creates clocks from external clock not locked
Status bit 2 : always logic '1'
Status bit 1 : LMK3200(2) PLL and clock distribution chip not locked to external clock
Status bit 0 : LMK3200(1) PLL and clock distribution chip not locked to external clock
If all these bits are not set then the operation of the firmware is unreliable

     -  adc calibration:

Calibration test result: Passed 5, Failed 0

If any modules fail calibration , check the clock status and open the FADC Align and Control browser page to rerun calibration for that module

   - white rabbit check:

     Base         Current     Difference
aida01 fault      0xc228 :      0xc274 :      76  
aida02 fault      0x5e98 :      0x5ee0 :      72  
aida03 fault      0x92d3 :      0x931f :      76  
aida04 fault      0x7a92 :      0x7ade :      76  
aida05 fault      0xedc0 :      0xedcc :      12  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

 

DAQ (WS2): everything ok

Online monitoring (WS 5): done;

  238   Tue Feb 22 16:09:08 2022 David, YuryCurrent status

Merger, Tape server (WS3) checks: DONE, writing ~1 1273kB/s, refreshing is fine

Leakage current (WS4): 3.46 uA , stable current

DAQ (WS2): temperature checked,

     - clock status:

Clock status test result: Passed 5, Failed 0

Understand status as follows
Status bit 3 : firmware PLL that creates clocks from external clock not locked
Status bit 2 : always logic '1'
Status bit 1 : LMK3200(2) PLL and clock distribution chip not locked to external clock
Status bit 0 : LMK3200(1) PLL and clock distribution chip not locked to external clock
If all these bits are not set then the operation of the firmware is unreliable

     -  adc calibration:

Calibration test result: Passed 5, Failed 0

If any modules fail calibration , check the clock status and open the FADC Align and Control browser page to rerun calibration for that module

   - white rabbit check:

         Base         Current     Difference
aida01 fault      0xc228 :      0xc272 :      74  
aida02 fault      0x5e98 :      0x5edf :      71  
aida03 fault      0x92d3 :      0x931d :      74  
aida04 fault      0x7a92 :      0x7adc :      74  
aida05 fault      0xedc0 :      0xedcc :      12  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

DAQ (WS2): everything ok

Online monitoring (WS 5): done

  237   Tue Feb 22 14:32:04 2022 David, YuryCurrent status

Merger, Tape server (WS3) checks: DONE, writing ~1 MB/s, refreshing is fine

Leakage current (WS4): 3.48 uA , stable current

DAQ (WS2): temperature checked,

     - clock status:
Clock status test result: Passed 5, Failed 0

Understand status as follows
Status bit 3 : firmware PLL that creates clocks from external clock not locked
Status bit 2 : always logic '1'
Status bit 1 : LMK3200(2) PLL and clock distribution chip not locked to external clock
Status bit 0 : LMK3200(1) PLL and clock distribution chip not locked to external clock
If all these bits are not set then the operation of the firmware is unreliable

     -  adc calibration:
Calibration test result: Passed 5, Failed 0

If any modules fail calibration , check the clock status and open the FADC Align and Control browser page to rerun calibration for that module

   - white rabbit check:

         Base         Current     Difference
aida01 fault      0xc228 :      0xc272 :      74  
aida02 fault      0x5e98 :      0x5edf :      71  
aida03 fault      0x92d3 :      0x931d :      74  
aida04 fault      0x7a92 :      0x7adc :      74  
aida05 fault      0xedc0 :      0xedcc :      12  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

DAQ (WS2): everything ok

Online monitoring (WS 5): done; first shots with something?

 

 

  236   Tue Feb 22 12:23:23 2022 CBSlow comparator thresholds lowered - test run

13:20 Stopped run.

Slow comp thresh 0x64 -> 0x32 on all FEEs

Checkload x2

 

New run - R55

Restarted Data Spy code

Zerod histograms

Data rate acceptable.

Dead channels seem to come back online

 

Please check for deadtime/idle time once we collect enough stats.

  235   Tue Feb 22 12:01:46 2022 RSS, OF, ML, CB Beam OFF, Target ON Run 54

01:00

Beam optimisation since 09:25.

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

Beam optimization will be done until 3 pm and will be then give for the experiment.

  234   Tue Feb 22 10:30:00 2022 RSS, OF, ML Beam OFF, Target ON Run 54

11:30

Beam optimisation. No beam since 09:25.

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

Beam optimization will be done until 3 pm and will be then give for the experiment.

  233   Tue Feb 22 10:06:50 2022 RSS, OF, MLPMT channel

YR00DA1ZB is the channel for PMT rate.

  232   Tue Feb 22 09:10:48 2022 OHAnalysis of R54_14
Results of the fortran analyser programs for R54_14

No timewarps observed in the data
No dead time observed either.
Lots of idle time but expected with high thresholds and low noise
  231   Tue Feb 22 08:31:25 2022 RSS, OF, ML Beam OFF, Target ON Run 54

09:30

Beam optimisation. No beam since 09:25.

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

FH and ML are optimising the beam.

  230   Tue Feb 22 07:27:28 2022 RSS, OF, ML Beam ON, Target ON Run 54

08:30

Beam ON ~1.0E7 pps

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

FH and ML are optimising the beam.

  229   Tue Feb 22 06:30:09 2022 RSS, OF Beam ON, Target ON Run 54

07:30

Beam ON ~8.0E6 pps

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

08:00 The beam is electron cooled by FH and the orbit of the beam is slighty changed.

  228   Tue Feb 22 05:47:18 2022 JM, CB, LV, PM, AKBeam ON, Target ON Run 54 - 2 events at 10MeV!

06:47

Beam ON ~2.0E6 pps, still fluctates between 1e6 and 6e6 (like a step function)

Target ON and open to the ring

YR09xxx ETarget ~1.8e-10 (saturated)

Target E4 inlet pressure 8.0E-9 mbar

Dump S3 ~ 6.6E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

  227   Tue Feb 22 05:12:38 2022 JM, CB, LV, PM, AKBeam ON, Target ON Run 54

06:00

Beam ON ~1.5E6 pps, still fluctates between 1e6 and 6e6 (like a step function)

Target ON and open to the ring

YR09xxx ETarget ~1.8e-10

Target E4 inlet pressure 8.2E-9 mbar

Dump S3 ~ 6.5E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

  226   Tue Feb 22 04:09:28 2022 JM, CB, LV, PM, AKBeam ON, Target ON Run 54 - first event at 10MeV !

05:00

Beam ON ~2E6 pps, still fluctates between 2e6 and 6e6 (like a step function)

Target ON and open to the ring

YR09xxx ETarget ~1.8e-10

Target E4 inlet pressure 8.3E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

  225   Tue Feb 22 03:05:06 2022 JM, CB, LV, PM, AKBeam ON, Target ON Run 54

04:00

Beam ON ~5.5E6 pps, but it fluctates between 2e6 and 6e6 (like a step function) in about each 5-10min

Target ON and open to the ring

YR09xxx ETarget ~1.8e-10

Target E4 inlet pressure 8.4E-9 mbar

Dump S3 ~ 7.0E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

ELOG V3.1.4-unknown